# Predicting Stellar Parameters of Massive Stars from Light Curves with Machine Learning

### v1

@@ 1 Introduction
Currently, fundamental stellar parameters such as $T_{\rm eff}$ and $g$ are acquired through labor-intensive spectroscopic observations. Notably, the IACOB project, which aimed to observe the largest possible number of massive stars, achieved just over 1000 stars in its first 12 years using the 2.56-m Nordic Optical Telescope, the 1.2-m Mercator telescope, and the 2.2-m ESO telescope (Simón-D´ıaz et al., 2011a; Simón-D´ıaz et al., 2015, 2020). In contrast, time-series photometry enables the monitoring of vastly larger samples of stars. For example, the Transiting Exoplanet Survey Satellite (TESS), launched in 2018, allows for the collection of light curve data from over 200,000 stars at various temporal resolutions ranging from 20 seconds to 30 minutes (Ricker et al., 2015). Likewise, the Vera C. Rubin Observatory’s Legacy Survey of Space and Time (LSST), which began operations in 2025, will repeatedly image the southern sky over a 10-year baseline, constructing light curves for millions of stars (Ivezić et al., 2019). Although these time-series photometric data provide rich temporal information, they do not directly yield key stellar parameters such as effective temperature, surface gravity, or spectroscopic luminosity. This growing imbalance underscores the importance of developing methods to extract stellar information directly from large-scale time-series photometry.
@@
Independently, numerous studies have demonstrated the predictive power of stellar light curves across a variety of contexts and datasets. For classification tasks, deep learning architectures have shown particular promise, with convolutional neural networks (CNNs) effectively capturing both spatial and temporal patterns in astronomical time-series data (Pasquet et al., 2019; Neira et al., 2020; Cui et al., 2024). Beyond classification, light curves have also proven valuable for regression tasks aimed at predicting stellar parameters. Both traditional algorithms and modern deep learning approaches have successfully extracted stellar parameters from light curve data of stars cooler than O and B types (Miller et al., 2015; Blancato et al., 2022). Motivated by the success of this prior work and by evidence that the red noise component of massive star light curves encodes stellar information, we investigate whether similar strategies can be applied to massive stars.
@@ 2.1 Data
The IACOB database currently contains high-resolution spectra of O and B-type stars from the HERMES (Raskin et al., 2011) and FIES (Telting et al., 2014) spectrographs. For O stars with at least one high-resolution spectrum, we used the stellar parameters derived following the methodologies outlined in Holgado et al. (2018). This approach involves using the iacob-broad (Simón-D´ıaz & Herrero, 2014) tool to measure line-broadening parameters, such as the projected rotational velocity $v$ sin $i$ and macroturbulence, through a combined Fourier transform and goodness-of-fit analysis. Subsequently, the iacob-gbat (Simón-D´ıaz et al., 2011b) tool and a grid of FASTWIND models (Puls et al., 2005) is employed to determine atmospheric parameters, including effective temperature, gravity, helium abundance, among others. A key strength of this methodology is that it ensures a homogeneous and consistent analysis across many O stars while retaining a semi-automatic component. Despite the high level of automation, a final visual inspection is required to validate the results and account for peculiarities that classic automated algorithms might overlook. After applying this full process, we obtained a sample of 257 O-type stars with detailed stellar parameter information from IACOB. The complete IACOB catalog will be available at https://research.iac.es/proyecto/iacob/iacobcat/.
We then combined TESS light curves with their corresponding stellar parameters derived from IACOB by cross-matching their TIC IDs, producing a dataset restricted to stars present in both datasets. After removing stars with any $\rm NaN$ flux values, our final regression dataset comprises 285 O star light curves from TESS with corresponding labeled stellar information from IACOB. Since TESS observes stars across multiple sectors, the 285 light curves correspond to 106 unique stars, with each sector treated as an independent observation. Each star is characterized by measured stellar parameters: effective temperature ($T_{\rm eff}$) and surface gravity ($\log_{10}g$), and a derived spectroscopic luminosity ($\log_{10}\mathscr{L}$) defined by Equation 1.
@@ 2.2 Machine learning approach
Figure 2: Scatter plots in both panels compare actual and predicted values for spectroscopic luminosity (top rows) and effective temperature (bottom rows). Panel (a) presents results from MLP1 and CNN1 models, while panel (b) presents results from MLP2 and CNN2 models. The red line in each plot indicates perfect agreement between predicted and actual values, serving as a reference for prediction accuracy.
@@ 3.1 MLP on fitted red noise parameters
We plot the performance of MLP1 and MLP2 in Figure 2. In the left panels of Figure 2a, MLP1 shows limited effectiveness overall. For predicting $\log_{10}\mathscr{L}$, the upper left panel yields $R^{2}=0.254_{-0.104}^{+0.188}$, indicating that predictions capture only about $25.4\%$ of the variance in the observed values, with poor alignment to the ideal prediction line (shown in red). Predictions of $T_{\rm eff}$ perform worse, with $R^{2}=0.123_{-0.218}^{+0.045}$ (bottom left panel). Predictions of $\log_{10}g$ achieve marginally better results, with $R^{2}=0.346_{-0.211}^{+0.051}$.
@@ 3.2 CNN on full periodogram
Both CNN1 and CNN2 trained on full periodogram inputs achieve superior performance in predicting $\log_{10}\mathscr{L}$ and $T_{\rm eff}$ compared to their respective MLP1 and MLP2 architectures that take the four fitted parameters $\alpha_{0}$, $\nu_{\rm char}$, $\gamma$, and $C_{\rm W}$ as input. As shown in the upper right panels of Figure 2a and Figure 2b, CNN1 and CNN2 achieve $R^{2}$ scores of $0.641_{-0.167}^{+0.107}$ and $0.664_{-0.120}^{+0.096}$, respectively, for predicting $\log_{10}\mathscr{L}$. These values represent a substantial improvement over the corresponding MLP models. Additionally, as shown in the lower right panels of Figure 2a and Figure 2b, CNN1 and CNN2 achieve $R^{2}$ scores of $0.443_{-0.234}^{+0.056}$ and $0.346_{-0.150}^{+0.117}$, respectively, for predicting $T_{\rm eff}$, which also outperforms their corresponding MLPs. CNN1 further achieves strong performance for predicting $\log_{10}g$, with an $R^{2}$ value of $0.771_{-0.195}^{+0.024}$, which also surpasses the best MLP1 results.
Notably, the superior performance of the CNN is not simply due to its larger parameter count. To confirm this, we tested various MLP architectures ranging from 6-89 million trainable parameters using the standardized periodogram as input and found that despite having comparable to significantly more trainable parameters, the MLP architectures still underperformed CNNs (see Tables A.3 and A.4). This indicates that not only is using the entire periodogram as input important, but also that the CNN architecture is specifically needed to extract the most information for stellar parameter prediction.
@@ 4 Discussion
Some methodological considerations merit further discussion. As shown in all the subfigures of Figure 2, our models tend to predict values clustering around the mean, possibly overlooking the nuances at both the low and high ends of the spectrum. This behavior is analogous to the class imbalance issue in classification tasks, where models are biased towards more frequently represented categories. Future work could explore strategies to enhance the variance in predictions, potentially improving the model’s ability to capture the full range of stellar parameter values. Additionally, the choice between using periodograms or light curves as input presents another important consideration. While periodograms transformed from light curves effectively capture periodic phenomena crucial for stellar characterization, they still represent a dimension reduction of the original light curve data. This transformation might lose potentially valuable information present in the raw time series. However, working directly with light curves presents its own challenges, including increased noise and the need for more complex input processing for machine learning models. The trade-offs between these approaches warrant further investigation as we seek to maximize the stellar information we can extract from time-series photometric data.
In conclusion, our work establishes that (1) O star light curves contain sufficient information to predict stellar parameters, and (2) CNNs trained on full periodograms recover substantially more stellar information than MLPs trained on fitted red noise parameters. As surveys like LSST collect unprecedented volumes of light curve data without accompanying spectroscopy, the ability to infer stellar parameters from light curves becomes increasingly valuable. Our approach enables the preliminary characterization of massive stars at scale and helping to prioritize targets for spectroscopic follow-up. Extending this methodology to include a broader range of stellar types will further enhance its utility in advancing our understanding of stellar physics in the era of large-scale surveys.

### v2

@@ 1 Introduction
Currently, fundamental stellar parameters such as $T_{\rm eff}$ and $g$ are acquired through labor-intensive spectroscopic observations. Notably, the IACOB project, which aimed to observe the largest possible number of massive stars, achieved just over 1000 stars in its first 12 years using the 2.56-m Nordic Optical Telescope, the 1.2-m Mercator telescope, and the 2.2-m ESO telescope (Simón-Díaz et al., 2011a; Simón-Díaz et al., 2015, 2020). In contrast, time-series photometry enables the monitoring of vastly larger samples of stars. For example, the Transiting Exoplanet Survey Satellite (TESS), launched in 2018, allows for the collection of light curve data from over 200,000 stars at various temporal resolutions ranging from 20 seconds to 30 minutes (Ricker et al., 2015). Likewise, the Vera C. Rubin Observatory’s Legacy Survey of Space and Time (LSST), which began operations in 2025, will repeatedly image the southern sky over a 10-year baseline, constructing light curves for millions of stars (Ivezić et al., 2019). Although these time-series photometric data provide rich temporal information, they do not directly yield key stellar parameters such as effective temperature, surface gravity, or spectroscopic luminosity. This growing imbalance underscores the importance of developing methods to extract stellar information directly from large-scale time-series photometry.
@@
Independently, numerous studies have demonstrated the predictive power of stellar light curves across a variety of contexts and datasets. For classification tasks, deep learning architectures have shown particular promise, with convolutional neural networks (CNNs) effectively capturing both spatial and temporal patterns in astronomical time-series data (Pasquet et al., 2019; Neira et al., 2020; Cui et al., 2024). Beyond classification, light curves have also proven valuable for regression tasks aimed at predicting stellar parameters. Both traditional algorithms and modern deep learning approaches have successfully extracted physical parameters such as rotation period and surface gravity directly from light curves in stars cooler than O and B stellar types (Miller et al., 2015; Blancato et al., 2022; kamai2025). Motivated by the success of this prior work and by evidence that the red noise component of massive star light curves encodes stellar information, we investigate whether similar strategies can be applied to massive stars.
@@ 2.1 Data
The IACOB database currently contains high-resolution spectra of O and B-type stars from the HERMES (Raskin et al., 2011) and FIES (Telting et al., 2014) spectrographs. For O stars with at least one high-resolution spectrum, we used the stellar parameters derived following the methodologies outlined in Holgado et al. (2018). This approach involves using the iacob-broad (Simón-Díaz & Herrero, 2014) tool to measure line-broadening parameters, such as the projected rotational velocity $v$ sin $i$ and macroturbulence, through a combined Fourier transform and goodness-of-fit analysis. Subsequently, the iacob-gbat (Simón-Díaz et al., 2011b) tool and a grid of FASTWIND models (Puls et al., 2005) is employed to determine atmospheric parameters, including effective temperature, gravity, helium abundance, among others. A key strength of this methodology is that it ensures a homogeneous and consistent analysis across many O stars while retaining a semi-automatic component. Despite the high level of automation, a final visual inspection is required to validate the results and account for peculiarities that classic automated algorithms might overlook. After applying this full process, we obtained a sample of 257 O-type stars with detailed stellar parameter information from IACOB. The complete IACOB catalog will be available at https://research.iac.es/proyecto/iacob/iacobcat/.
We then combined TESS light curves with their corresponding stellar parameters derived from IACOB by cross-matching their TIC IDs, producing a dataset restricted to stars present in both datasets. After removing stars with any $\rm NaN$ flux values, our final regression dataset comprises 285 O star light curves from TESS with corresponding labeled stellar information from IACOB. Since TESS observes stars across multiple sectors, the 285 light curves correspond to 106 unique stars, with each sector treated as an independent observation. The 106 unique O stars used comprise 19 early ($\leq\rm O6$), 53 mid ($\rm O6<x\leq O8.5$), and 34 late ($\geq\rm O9$) type stars, with luminosity classes distributed as 53 dwarfs (IV – V), 16 giants (III), and 37 supergiants (I – II). Each star is characterized by measured stellar parameters: effective temperature ($T_{\rm eff}$) and surface gravity ($\log_{10}g$), and a derived spectroscopic luminosity ($\log_{10}\mathscr{L}$) defined by Equation 1.
2.2 Machine learning approach
@@
Figure 2: Scatter plots in both panels compare actual and predicted values for spectroscopic luminosity (top rows), effective temperature (second rows), and surface gravity for MLP1/CNN1 (bottom row of Panel (a)). Panel (a) presents results from MLP1 and CNN1 models, while panel (b) presents results from MLP2 and CNN2 models. The red line in each plot indicates perfect agreement between predicted and actual values, serving as a reference for prediction accuracy.
@@ 3.1 MLP on fitted red noise parameters
We plot the performance of MLP1 and MLP2 in Figure 2. In the left panels of Figure 2a, MLP1 shows limited effectiveness overall. For predicting $\log_{10}\mathscr{L}$, the upper left panel yields $R^{2}=0.254_{-0.104}^{+0.188}$, indicating that predictions capture only about $25.4\%$ of the variance in the observed values, with poor alignment to the ideal prediction line (shown in red). Predictions of $T_{\rm eff}$ perform worse, with $R^{2}=0.123_{-0.218}^{+0.045}$ (middle left panel). Predictions of $\log_{10}g$ achieve marginally better results, with $R^{2}=0.346_{-0.211}^{+0.051}$ (bottom left panel).
@@ 3.2 CNN on full periodogram
Both CNN1 and CNN2 trained on full periodogram inputs achieve superior performance in predicting $\log_{10}\mathscr{L}$ and $T_{\rm eff}$ compared to their respective MLP1 and MLP2 architectures that take the four fitted parameters $\alpha_{0}$, $\nu_{\rm char}$, $\gamma$, and $C_{\rm W}$ as input. As shown in the upper right panels of Figure 2a and Figure 2b, CNN1 and CNN2 achieve $R^{2}$ scores of $0.641_{-0.167}^{+0.107}$ and $0.664_{-0.120}^{+0.096}$, respectively, for predicting $\log_{10}\mathscr{L}$. These values represent a substantial improvement over the corresponding MLP models. Additionally, as shown in the middle right panel of Figure 2a and lower right panel of Figure 2b, CNN1 and CNN2 achieve $R^{2}$ scores of $0.443_{-0.234}^{+0.056}$ and $0.346_{-0.150}^{+0.117}$, respectively, for predicting $T_{\rm eff}$, which also outperforms their corresponding MLPs. CNN1 further achieves strong performance for predicting $\log_{10}g$ (lower right panel of Figure 2a), with an $R^{2}$ value of $0.771_{-0.195}^{+0.024}$, which also surpasses the best MLP1 results.
[…]
@@ 4 Discussion
Some methodological considerations merit further discussion. As shown in all the subfigures of Figure 2, our models tend to predict values clustering around the mean, possibly overlooking the nuances at both the low and high ends of the spectrum. This pattern is consistent with the non-uniform composition of our sample reported in Section 2.1 (19 early, 53 mid, 34 late; 53 dwarfs, 16 giants, 37 supergiants), which yields fewer training examples at more extremes. This behavior is analogous to the class imbalance issue in classification tasks, where models are biased towards more frequently represented categories. Future work could explore strategies to enhance the variance in predictions via methods such as loss re-weighting, stratified train/test splits, and inverse-frequency sampling, potentially improving the model’s ability to capture the full range of stellar parameter values.
Additionally, the choice between using periodograms or light curves as input presents another important consideration. While periodograms transformed from light curves effectively capture periodic phenomena crucial for stellar characterization, they still represent a dimension reduction of the original light curve data. This transformation might lose potentially valuable information present in the raw time series. However, working directly with light curves presents its own challenges, including increased noise and the need for more complex input processing for machine learning models. The trade-offs between these approaches warrant further investigation as we seek to maximize the stellar information we can extract from time-series photometric data.
[…]
In conclusion, our work establishes that (1) O star light curves contain sufficient information to predict stellar parameters, and (2) CNNs trained on full periodograms recover substantially more stellar information than MLPs trained on fitted red noise parameters. As surveys like LSST collect unprecedented volumes of light curve data without accompanying spectroscopy, the ability to infer stellar parameters from light curves becomes increasingly valuable. Our approach enables the preliminary characterization of massive stars at scale and helping to prioritize targets for spectroscopic follow-up. Extending this methodology to systematically include a broader range of stellar types will further enhance its utility in advancing our understanding of stellar physics in the era of large-scale surveys.